# Generalization of Task Parameterized Dynamical Systems using Gaussian Process Transportation

### v1

@@ Abstract
Learning from Interactive Demonstrations has revolutionized the way non-expert humans teach robots. It is enough to kinesthetically move the robot around to teach pick-and-place, dressing, or cleaning policies. However, the main challenge is correctly generalizing to novel situations, e.g., different surfaces to clean or different arm postures to dress. This article proposes a novel task parameterization and generalization to transport the original robot policy, i.e., position, velocity, orientation, and stiffness. Unlike the state of the art, only a set of points are tracked during the demonstration and the execution, e.g., a point cloud of the surface to clean. We then propose to fit a non-linear transformation that would deform the space and then the original policy using the paired source and target point sets. The use of function approximators like Gaussian Processes allows us to generalize, or transport, the policy from every space location while estimating the uncertainty of the resulting policy due to the limited points in the task parameterization point set and the reduced number of demonstrations. We compare the algorithm’s performance with state-of-the-art task parameterization alternatives and analyze the effect of different function approximators. We also validated the algorithm on robot manipulation tasks, i.e., different posture arm dressing, different location product reshelving, and different shape surface cleaning. A video of the experiments can be found here: https://youtu.be/FDmWF7K15KU.
@@ I Introduction
By enhancing the policy with task parameterization [2], robots can generalize their learned knowledge to different variations of the same task, thus promoting scalability and data efficiency in robot learning, allowing robots to learn faster and adapt to new scenarios. For instance, a robot can be trained to clean surfaces with a reduced set of shapes, to dress an arm in a certain configuration, or to pick objects with a certain shape and place them on the right shelf. Ideally, the robot would generalize the learned skill to novel situations without extensive retraining. This paper proposes a way of transfer, or “transport”, the original learned behavior from the old to the new situation.
For example, let us imagine teaching a robot how to draw on a flat canvas, as depicted in Fig. 1. The robot can learn to imitate the human on the same flat canvas, however, when the desired surface to reproduce the task is changed, i.e., the blue curvy surface, the user may have to teach the task again on the new surface. Instead, by knowing the correspondence of a set of points from the two surfaces, we propose to learn a function that could generalize the policy from the original parameter distribution to the new one.
The learned function is locally deforming the space according to the new location of the tracked points and this can be used to generalize the learned velocity field, learned stiffness and orientation. Fig. 2, depicts how the trajectory and the learned desired dynamics of drawing a letter C depicted as red dots, on a flat surface, depicted in green; the robot would learn the desired behavior, i.e., the velocity depicted as black arrows, form any position of the space. However, when the desired surface to reproduce the task is changed, i.e., the blue curvy surface, the user may have to teach the task again on the new surface. Instead, by knowing the correspondence of a set of points from the two surfaces, we show how to learn a function that could generalize the policy from the original surface shape to the new one.
@@
Figure 1: Example of Policy Transportation. The human demonstrates to a robot how to perform a task on a flat canvas. Then, the robot, when facing a new curved canvas, “transports” its knowledge in the new situation by adapting the end effector velocity, orientation, and stiffness to correctly adapt the drawing on the new canvas.
@@
Our algorithm contributes with the formalization and testing of a policy transportation theory that can
@@
  • transport the demonstration from the original space to the new space, see Sec. IV-B;
@@
  • transport the velocity field, end-effector orientations, stiffness and damping by exploiting the derivative of the transportation mapping, see Sec. IV-C;
  • estimates the final uncertainty due to the reduced set of demonstrations and the estimated uncertainty in the transportation map, see Sec. IV-E.
@@
The same algorithm was tested on generalizing complex manipulation tasks like cleaning surfaces with different shapes, picking and placing objects at other locations, and dressing an arm in various configurations, see Sec. VI.
Figure 2: The demonstrations (in red) are given on the green surface, and the learned dynamics, depicted as arrows, are learned from them. Later, the demonstration and the dynamics are projected on another curved blue surface using the proposed Gaussian Process Transportation.
@@ II Related Works
An approach to adapt the DMPs via points is addressed in [4], but it demands combining several DMPs for a single task. Alternatively, a roto-translation can be applied to the original dynamical systems according to the tracked frame or points in the environment.
@@
Approaches for modeling and generalizing demonstrations that have shown improved performances with respect to the DMPs are Probabilistic Movement Primitives (ProMPs) [5]. ProMPs model the distribution over the demonstrations that capture temporal correlation and correlations between the DoFs using a linear combination of weights and a set of manually designed basis functions. Adaptation to new task parameters or via points is achieved using Gaussian conditioning. While this approach allows modeling the structure and variance of the observed data in the absolute reference frame, the generalization to the new task parameters is satisfactory only within the confidence bound of the demonstration data. For example, showing many demonstrations for different goal points, the probabilistic model can be conditioned on a novel object position and retrieve the most probable trajectory that brings the robot to that final position.
However, when learning reactive policies, i.e., a function of the state and not of the phase of the motion, the use of ProMPs is limited since the number of basis functions overgrows with the dimension of the input, limiting its applications.
Kernelized Movement Primitives (KMP) [6] proposed a non-parametric skill learning formulation.
@@
This formulation allows modulation of the recorded trajectories to new via points, obtaining the deformation of the original movement primitives given the temporal correlation of the demonstration and the via point, calculated with the kernel function. However, the user must specify the time and the corresponding waypoint to deform the original trajectory or rely on a heuristic that, for example, matches each waypoint with the closest point in the trajectory.
Gaussian Mixture Models (GMM) [7, 8, 9] have successfully been employed in modeling demonstration, endowing with a successful generalization in its task parameterized version (TP-GMM) [2].
Given a set of reference frames that are tracked during demonstration and execution, the central idea of TP-GMM is the local projection of the demonstrations in each of the local reference frames and encoding each model as a mixture of Gaussians [10, 2, 11, 12]. The local models are then fused in global coordinates, using the Product of Gaussian (PoG), and a new motion is rolled out from the resulting mixture model. This approach, however, requires many demonstrations to fit the model and does not scale well with the increasing number of task frames. This is because the PoG does not scale well when dealing with many reference frames and can lead to undesirable generalizations.
@@
The generalization of the demonstrations, encoded as a chain of events in a graph, with respect to multiple via points, can be done using Laplacian Editing [13, 14]. It uses Laplace-Betromi operator, a well-known algorithm in the computer graphics community, to deform meshes [15], to encode geometric trajectory properties and generate deformed trajectories using task constraints, i.e., new via points. The operator ensures a smooth deformation of the trajectory through the via points.
However, this approach is very specific for trajectory reshaping and requires explicit knowledge of the new via-point for some trajectory nodes.
In this paper, we relax the need to explicitly specify the new via-points for a specific node in the trajectory since the via-point definition can be error-prone and requires ad-hoc algorithms.
Moreover, while all the previous approaches only address generalizing a demonstration/robot trajectory, our proposed approach relaxes the requirement of generalizing only trajectories. It provides the means also to generalize the velocity field of the original dynamics, as well as the orientation and the stiffness.
For generalization, we use a combination of linear transformation and non-linear deformation to transport demonstrations to new situations while estimating the process uncertainty. We validated the algorithm on three use cases: robot reshelving, robot dressing, and robot cleaning.
@@
Robot Reshelving Authors of [16] propose, within the realm of robotic retail automation, to enable non-expert supermarket employees to teach a robot a reshelving task and then adeptly generalize its learned policy to accommodate diverse task situations. The generalization of the policy for varying object locations is achieved by switching between the dynamical system learned between the object and the goal frame. However, the switching strategy entails having a good prior on when to switch and all the possible implications of generating instability by suddenly changing the policy online. TP-GMM alternatives [2], solved the problem of the switching by obtaining the final GMM as the product of the relative models; however more than one demonstration is necessary to fit an informative model.
@@
Robot Dressing Robot dressing is a challenging task since it includes manipulation of deformable objects, and the margin of error to correctly go through the human arm is very low.
Task parameterized dynamical system has been applied to learn the dressing task in the robotics research. The dressing demonstrations w.r.t. the wrist and the shoulder of a human arm have been used to learn a dressing policy via DMP [17], TP-HMM [18] and a TP-GMM [19].
Robot Surface Cleaning Efficient and fast generalization of robotic surface cleaning is achieved using task-parameterized learning.
@@
In [20] the cleaning dynamics is the sum of two dynamical systems, one that learns the desired motion on the surface and another that computes the modulation term on the desired force to apply on the perpendicular direction of the surface (where the shape is known a priori). This second term is learned as a non-linear function that allows learning larger forces in a region of the surface compared to others.
The shape of the surface can also be estimated using the wrench measured with a force-torque sensor attached at the end-effector; for instance, [21] generalizes the polishing task on the novel curved surface by adapting the orientation and the direction of the contact force such that to minimize perceived torque.
The following section will provide some background on the main concepts necessary for learning a policy from demonstration and generalize it using Gaussian Process Transportation.
III Preliminaries
We review the learning of policies from demonstration using a dynamical system formulation in Sec. III-A and the specific use of a Gaussian Process (GP) for learning the policy itself or the transportation policy in Sec. III-B.
III-A Interactive Learning of Dynamical System from Demonstrations
A non-linear dynamical system (DS) can be described by
$\dot{\bm{x}}=f(\bm{x})$ (1)
where $\bm{x}$ is the state of the system, such as the end effector or the joint position, while $\bm{\dot{x}}$ is the observed state transition.
The demonstrated dynamics are captured by a non-linear function $f$ that is agnostic to the function approximation used. However, the resulting properties for out-of-distribution predictions depend on it. For example, the robot may extrapolate the wrong behavior and make the motion to diverge in undesired regions of the workspace. To enforce stability of the motion dynamics, the Stable Estimator of Dynamical Systems (SEDS) [22], constrains the properties of each of the linear models of a Gaussian mixture to enforce global asymptotic stability. Alternatively, the global dynamics can be coupled with latent (linear) stable dynamics, using a deep neural network projector to ensure the global stability of the learned task dynamics [23, 24].
Differently, [25] proposes to minimize the epistemic uncertainties predicted as an output of a GP dynamical system to attract the motion in regions with high confidence, i.e., close to the training set. Moreover, [25] also proposes a way of updating the learned attractor and stiffness field online, to learn for example, how to apply enough pressure on a table, and clean it successfully. Fig. 2 shows the dynamics learned from the drawing of a letter C from multiple demonstrations, where the attractor delta in z is set to be constantly negative, i.e., the vector arrows point under the surface. Since, in the context of this article, the policy and the transportation are learned with a Gaussian Process, the next section will introduce the fundamental equations used to fit and predict mean and variance using a GP (or its approximation).
III-B Gaussian Process Regression
where $\bm{X}_{*}\in\mathbb{R}^{n\times m}$ are the evaluation inputs and $\bm{X}\in\mathbb{R}^{N\times m}$ and $\bm{y}\in\mathbb{R}^{N\times l}$ are the training inputs and outputs; the correlations $\bm{K}_{\bm{X},\bm{X}}\in\mathbb{R}^{N\times N}$, $\bm{K}_{\bm{X}_{*},\bm{X}}\in\mathbb{R}^{n\times N}$, $\bm{K}_{\bm{X}_{*},\bm{X}_{*}}\in\mathbb{R}^{n\times n}$ are the correlation matrix between every input training point, between testing and training points and between testing points.
Every entry of this matrices is computed using a kernel function, for example, the squared exponential kernel, defined as
where $\bm{x_{i}},\bm{x_{j}}$ are any pair of training-training, testing-training, or testing-testing data points; $\sigma_{p}$ and $\ell$ are the kernel hyperparameters and, together with the likelihood noise $\sigma_{n}$ are optimized with a maximization of the log-likelihood of the given data [26].
Given the computationally expensive inversion of the covariance matrix that grows with the number of data points, optimizing kernel hyperparameters and inference becomes prohibitive when having big datasets. Therefore, variational inference aims to approximate the true posterior $p(\bm{f}|\bm{X},\bm{y})$ with a simpler distribution $q(\bm{f})$. Stochastic Variational Gaussian Process (SVGP) optimizes the location and values of a set of pseudo data, usually referred to as inducing points $\bm{Z}$ and a variational distribution $q(\bm{u})$. The parameters of the kernel and the variational parameter of the approximated distribution are optimized using the evidence lower bound (ELBO) as the evidence of the data[27].
The use of SVGP, in the context of this paper, can make the computation significantly faster when fitting a transportation function with many points from the source and target distribution; however, any of the proposed algorithms, formalized in the next section, is independent of the approximation nature of the GP.
IV Policy Transportation
We learn and correct a manipulation policy from an interactive demonstration [1], using, for example, a GP as the function approximator to fit the demonstration dynamics.
However, although the task execution would succeed from any given starting configuration of the robot, it will fail to generalize if the task is changed, e.g., if the object to pick is moved or if the robot faces a differently shaped surface to clean.
Intending to find a task parameterization that scales from pick-and-place to continuous surface, we assume to track a set of environment-specific points that are descriptive of the situation. For example, when picking a box, the eight corners are tracked, or when cleaning a surface, a point cloud representation is used. This is the most straightforward yet most general task parameterization while being technically feasible, given the current development of LiDAR and depth camera technology.
We define the tracked $N$ point recorded in the demonstration scenario as the source distribution, i.e.,
$\bm{\mathcal{S}}=\{(x_{s,i},y_{s,i},z_{s,i})\}_{i=1}^{N}$ (4)
while the moved points in the new scenario are defined as the target distribution, i.e.,
$\bm{\mathcal{T}}=\{(x_{t,i},y_{t,i},z_{t,i})\}_{i=1}^{N}.$ (5)
Figure 3: 2D transportation. Distribution match depicts the source and the target distribution correspondence used to train the transportation function. Source Distribution depicts a grid of points in the original space. Linear Transformation shows the effect on the original grid when only a linear transformation is used to match source and target distribution. GP Transportation captures the deformation of the space when the source points are forced to match the target ones.
We assume that the points of target and source distribution are already paired. Many algorithms are available to (optimally) pair the two distributions [28]; hence this is not the focus of our work.
We define a map $\bm{\phi}$ such that each point $s_{i}$ in $S$ is paired with one and only one point $t_{j}$ in $T$. This can be represented as:
$\bm{\phi}:\bm{\mathcal{S}}$ $\to\bm{\mathcal{T}}$ (6)
$s_{i}$ $\mapsto t_{j}$
where $i,j\in\{1,2,\ldots,n\}$. We aim to find the function that maps from the source space to the target space, given the evidence of the input-output pairs from the source and target distribution. Estimating a continuous process allows the deformation of the complete space to match the source and the target distribution, as depicted in Fig. 3.
The structure of the function $\bm{\phi}$ that we want to approximate can be any nonlinear function that maps any point of the Cartesian space to itself, e.g., $\bm{\phi}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$. However, in the context of this article, we consider the transportation function to have the following definition,
$\bm{\phi}(\bm{x}):=\bm{\gamma}(\bm{x})+\bm{\psi}(\bm{\gamma}(\bm{x}))$ (7)

### v2

# Generalizable Motion Policies through Keypoint Parameterization and Transportation Maps

@@ Abstract
Learning from Interactive Demonstrations has revolutionized the way non-expert humans teach robots. It is enough to kinesthetically move the robot around to teach pick-and-place, dressing, or cleaning policies. However, the main challenge is correctly generalizing to novel situations, e.g., different surfaces to clean or different arm postures to dress. This article proposes a novel task parameterization and generalization to transport the original robot policy, i.e., position, velocity, orientation, and stiffness. Unlike the state of the art, only a set of keypoints is tracked during the demonstration and the execution, e.g., a point cloud of the surface to clean. We then propose to fit a nonlinear transformation that would deform the space and then the original policy using the paired source and target point sets. The use of function approximators like Gaussian Processes allows us to generalize, or transport, the policy from every space location while estimating the uncertainty of the resulting policy due to the limited task keypoints and the reduced number of demonstrations. We compare the algorithm’s performance with state-of-the-art task parameterization alternatives and analyze the effect of different function approximators. We also validated the algorithm on robot manipulation tasks, i.e., different posture arm dressing, different location product reshelving, and different shape surface cleaning. A video of the experiments can be found here: https://youtu.be/bE6uOnAQBLo.
@@ I Introduction
By enhancing the policy with task parameterization [2], robots can generalize their learned knowledge to different variations of the same task, thus promoting scalability and data efficiency in robot learning, allowing robots to learn faster and adapt to new scenarios. For instance, a robot can be trained to clean surfaces with a reduced set of shapes, to dress an arm in a certain configuration, or to pick objects with a certain shape and place them on the right shelf. A versatile task parametrization to describe different situations, such as object pose, arm posture, and surface shape, is through a set of keypoints [3].
However, the current state of the art in task parameterization struggles to scale when managing numerous task parameters, such as describing a surface as a set of points, and fails to generalize effectively for policy states that are not in close proximity to the parameterization points due to the covariate shift.
We propose the learning of a “transportation” map that is trained to align a set of keypoints from the original task configuration (e.g., a flat surface) to a new configuration (e.g., a curved surface) as illustrated in Fig. 1. This map is then used to generalize the labels of the original policy, such as position, velocity, stiffness, and orientation. Given that the original policy’s set of labels might have been interactively adjusted or aggregated based on human feedback [1, 4], transforming the demonstration alone is insufficient. This limitation makes trajectory reshaping methods inappropriate for achieving the generalization goal [5].
@@
Figure 1: Example of Policy Transportation. The human demonstrates to the robot how to perform a task on a flat canvas. Then, the robot, when facing a new curved canvas, “transports” its knowledge in the new situation by adapting the end effector velocity, orientation, and stiffness to correctly adapt the drawing on the new canvas.
Therefore, we propose a method that not only works for the generalization of demonstrations encoded as trajectories but also if they are encoded as a set of independent labels. The proposed algorithm will “transport” the original labels on the new task configuration where a new policy can be re-fitted, making the generalization algorithm agnostic to the method used for policy learning, e.g., DMP [6], KMP [7], GP [4], NN [8], GMM [9], LPV [10], etc.
Moreover, our goal is to generalize out-of-distribution policies, e.g., motions that are far away from the table to clean. For this reason, it is important to define a transportation function that has desired out-of-distribution (o.o.d.) properties and is able to correctly quantify the uncertainties of the generalization. Ideally, the robot would generalize the learned skill to novel o.o.d. situations without requiring the aggregation of hundreds of thousands of demonstrations [11], regulate its stiffness as a function of the uncertainty [12], or actively ask for data in uncertain situations [4].
@@
The next sections will highlight how this paper contributes to the field of robot policy generalization with the formalization and testing of a policy transportation theory, Secs. III-A-III-D that can
@@
  • transport the demonstration position labels from the original task space to the new (deformed) space, see Sec. III-E;
@@
  • transport the position labels, velocity labels, end-effector orientations, stiffness, and damping by exploiting the derivative of the transportation mapping, see Sec. III-F;
  • estimate the final uncertainty due to the reduced set of demonstrations and the estimated uncertainty in the transportation map, see Sec. III-H.
@@
The proposed algorithm was tested on generalizing complex manipulation tasks like cleaning surfaces with different shapes, picking and placing objects at other locations, and dressing an arm in various configurations, see Sec. V.
@@ II Related Works
An approach to adapt the DMPs via points is addressed in [13], but it demands combining several DMPs for a single task. Alternatively, a roto-translation can be applied to the original dynamical systems according to the tracked frame or points in the environment and generalize the task with respect to the goal [14]. Approaches for modeling and generalizing demonstrations that have shown improved performances with respect to the DMPs are Probabilistic Movement Primitives (ProMPs) [15]. ProMPs model the distribution over the demonstrations that capture temporal correlation and correlations between the DoFs using a linear combination of weights and a set of manually designed basis functions. Adaptation to new task parameters or via points is achieved using Gaussian conditioning. While this approach allows modeling the structure and variance of the observed data in the absolute reference frame, the generalization to the new task parameters is satisfactory only within the confidence bound of the demonstration data. For example, showing many demonstrations for different goal points, the probabilistic model can be conditioned on a novel object position and retrieve the most probable trajectory that brings the robot to that final position. However, when learning reactive policies, i.e., a function of the state and not of the phase of the motion, the use of ProMPs is limited since the number of basis functions overgrows with the dimension of the input, limiting its applications.
@@
Kernelized Movement Primitives (KMP) [7] proposed a non-parametric skill learning formulation. This formulation allows modulation of the recorded trajectories to new via points, obtaining the deformation of the original movement primitives given the temporal correlation of the demonstration and the via point, calculated with the kernel function. However, the user must specify the time and the corresponding waypoint to deform the original trajectory or rely on a heuristic that, for example, matches each waypoint with the closest point in the trajectory.
@@
The generalization of the demonstrations, encoded as a chain of events in a graph, with respect to multiple via points, can be done using Laplacian Editing [16, 5]. It uses the Laplace-Beltrami operator, a well-known algorithm in the computer graphics community, to deform meshes [17], to encode geometric trajectory properties and generate deformed trajectories using task constraints, i.e., new via points. The operator ensures a smooth deformation of the trajectory through the via points. Similarly, [18] proposed a conditioning operation to specify the desired waypoint at a specific time of the motion. However, this approach is very specific for trajectory reshaping or movements that are encoded as a function of time and requires explicit knowledge of the new via point for some trajectory nodes or time instances.
Gaussian Mixture Models (GMM) [19, 20, 21] have successfully been employed in modeling demonstration, endowing with a successful generalization in its task parameterized version (TP-GMM) [2]. Given a set of reference frames that are tracked during demonstration and execution, the central idea of TP-GMM is the local projection of the demonstrations in each of the local reference frames and encoding each model as a mixture of Gaussians [22, 2, 23, 24]. The local models are then fused in global coordinates, using the Product of Gaussians (PoG), and a new motion is rolled out from the resulting mixture model. This approach, however, requires many demonstrations to fit the model and does not scale well with the increasing number of task frames. This is because the PoG does not scale well when dealing with many reference frames and can lead to undesirable generalizations. Moreover, [25] also uses a Gaussian Mixture Regression to adopt a linear limit cycle to a nonlinear one by learning a modification of the motion amplitude as a function of the phase. However, this is very specific to phase-dependent motions, where the learned phase-dependent map would never have to be evaluated for out-of-distribution phases.
A different task-parameterized approach to generalization as formulated in [26] uses Task-Parameterized Equation Learner Networks that have as features the task parameters and the time of each demonstration and given a set of demonstrations it can regress the equations that describe the desired position as a function of time and desired task parameters, such as the desire hight of the goal. By avoiding overfitting, the method also performs well on task parameters that are outside the boundary of the observed range. However, since the parameters are considered as extra features of the regression, very far extrapolations are not possible; moreover, the method requires multiple time-dependent demonstrations to regress the equations. We aim to relax both of these requirements and to extrapolate to extreme task parameter changes.
In this paper, we rely on a convenient 3D keypoint parametrization since the points could be extracted from a tracked object/reference frame or a point cloud. Moreover, we relax the need to explicitly specify the new via points for a specific node in the trajectory since the via point definition, i.e., which point of the trajectory should be moved, can be error-prone and require ad-hoc assignment algorithms between the keypoints and the trajectory. Yet, using keypoints is a double-edged sword. Although keypoints are a more expressive description of the task [3], if they are not moving rigidly, we cannot simply roto-translate our policy [27, 28]. We would have to rely on some planning [29] or sim-2-real RL [30]. We are the first to contribute a nonlinear transformation in this context of imitation learning to modify the original labels and show successful zero-shot motion generalization to different surface shapes, different arm configurations in dressing, and different object locations and goal locations.
The requirement of having a bijective map is usually used to enforce the asymptotic stability of nonlinear policy [31] [32] or to couple the motion of two manipulators to perform bilateral teleoperation [33]. However, in this article, we relax the requirement on the invertibility of the map, similarly to [34], and moreover, we use the map to transfer dynamics, orientation, and stiffness, and not only to couple the motion of two manipulators [33].
@@
Robot Reshelving: Authors of [35] propose, within the realm of robotic retail automation, to enable non-expert supermarket employees to teach a robot a reshelving task and then adeptly generalize its learned policy to accommodate diverse task situations. The generalization of the policy for varying object locations is achieved by switching between the dynamical system learned between the object and the goal frame. However, the switching strategy entails having a good prior on when to switch and all the possible implications of generating instability by suddenly changing the policy online. TP-GMM alternatives [2], solved the problem of the switching by obtaining the final GMM as the product of the relative models; however more than one demonstration is necessary to fit an informative model.
@@
Robot Dressing: Robot dressing is a challenging task since it includes manipulation of deformable objects, and the margin of error to correctly go through the human arm is very low. Task parameterized dynamical systems have been applied to learn dressing tasks; for example, the dressing demonstrations w.r.t. the wrist and the shoulder of a human arm have been used to learn a dressing policy via DMP [36], TP-HMM [37] and a TP-GMM [38].
@@
Robot Surface Cleaning: Efficient and fast generalization of robotic surface cleaning has been achieved using task-parameterized learning. In [39], the cleaning dynamics is the sum of two dynamical systems, one that learns the desired motion on the surface and another that computes the modulation term on the desired force to apply on the perpendicular direction of the surface (where the shape is known a priori). This second term is learned as a nonlinear function that allows learning larger forces in a region of the surface compared to others. The shape of the surface can also be estimated using the wrench measured with a force-torque sensor attached at the end-effector; for instance, [40] generalizes the polishing task on the novel curved surface by adapting the orientation and the direction of the contact force to minimize the perceived torque.
The following section formalizes the concept of policy generalization using transportation maps.
III Policy Transportation
III-A Problem Setting
When learning a robot policy from (interactive) demonstration, a set of desired state-action pairs is aggregated. Throughout the paper, we denote the available policy labels as:
  • The set of $M$ Cartesian position labels $\bm{\mathcal{X}}=\{\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{M}\}$, where $\bm{x}_{i}\in\mathbb{R}^{3}$.
  • The set of $M$ Cartesian velocity labels $\bm{\mathcal{\dot{X}}}=\{\dot{\bm{x}}_{1},\dot{\bm{x}}_{2},\ldots,\dot{\bm{x}}_{M}\}$, where $\dot{\bm{x}}_{i}\in\mathbb{R}^{3}$.
[…]
  • The set of $M$ Cartesian stiffness labels $\bm{\mathcal{K}}=\{\bm{K}_{1},\bm{K}_{2},\ldots,\bm{K}_{M}\}$, where $\bm{K}_{i}\in S^{+}_{3}$.
  • The set of $M$ Cartesian damping labels $\bm{\mathcal{D}}=\{\bm{D}_{1},\bm{D}_{2},\ldots,\bm{D}_{M}\}$, where $\bm{D}_{i}\in S^{+}_{3}$.
where $\mathbb{R}^{3}$ denotes the 3-dimensional Euclidean space, $SO(3)$ represents the special orthogonal group of 3x3 rotation matrices, and $S^{+}_{3}$ is the manifold of $\times 3$ 3 symmetric positive semi-definite matrices.
To develop a task parameterization that scales from pick-and-place to continuous surfaces, we track a set of environment-specific keypoints that describe the situation. Our task parameterization comprises $N$ tracked points in $\mathbb{R}^{3}$, recorded in the demonstration scenario as the source distribution, while the corresponding moved points in the new scenario are defined as the target distribution, defined as two ordered sets
  • $\bm{\mathcal{S}}=\{\bm{s}_{1},\bm{s}_{2},\ldots,\bm{s}_{N}\}$, where $\bm{s}_{i}\in\mathbb{R}^{3}$.
  • $\bm{\mathcal{T}}=\{\bm{t}_{1},\bm{t}_{2},\ldots,\bm{t}_{N}\}$, where $\bm{t}_{i}\in\mathbb{R}^{3}$.
Here, we ask these sets to be ordered as we assume that the points of the target and source distributions are already paired. This is a reasonable assumption as this pairing can either be done manually or autonomously via state-of-the-art solutions [41, 42].
[…]
III-B Proposed solution: Policy transportation
We propose to solve this challenge via a two step process (see Fig. 5)
  • First, we learn a policy transportation, mapping the labels $(\bm{\mathcal{X}},\bm{\mathcal{\dot{X}}})$ in the target space $(\bm{\mathcal{\hat{X}}},\bm{\mathcal{\dot{\hat{X}}}})$ in a way coherent with the keypoints;
  • Second, we use state-of-the-art techniques to learn the dynamics $g$ from $(\bm{\mathcal{\hat{X}}},\bm{\mathcal{\dot{\hat{X}}}})$.
The rest of the section will, therefore, focus on defining, analyzing, and learning the transportation map of step 1. We start here by introducing the general definition of the policy transportation. If the number of keypoints is larger than the size of the embedding space (3 in this work), it is, in general, not possible to match them with a linear transformation. Therefore, our policy transportation will necessarily need to be nonlinear. Let us define it as a continuously differentiable ($C^{1}$) map
$\bm{\phi}:\mathbb{R}^{3}\to\mathbb{R}^{3}.$ (1)
Note that our secondary goal can also be achieved as a beneficial by-product of the proposed solutions. We will discuss later in this section how all the remaining labels will be transformed and the uncertainty quantified.
III-C Desirable properties for ${\phi}$
Desired properties: The following are desirable properties $\phi$ of(1) to fulfill
  • (i)
    all keypoints are properly matched - i.e., $\bm{\mathcal{T}}=\bm{\phi}(\bm{\mathcal{S}})$,
  • (ii)
    the policy transportation is a change of coordinates - i.e., $\det(J_{\phi})\neq 0$ everywhere.
The reason behind the first property should be clear, as it directly reflects the problem statement.
The reason for considering the second property is that, in general, we would like $\phi$ to transform the labels in such a way it conserves stability properties. This is not necessary to our method, but it would be beneficial since conserving stability properties may result in better performance. Assume that $(\bm{\dot{\mathcal{X}}},\bm{\mathcal{X}})$ are samples extracted from an ideal underlying dynamical system $\dot{x}=f(x)$. Then, we would like the policy $g$ to have the same stability properties as $f$. For example, if $f$ is a contractive field, then the dynamics in the new coordinates should also be contractive. A necessary and sufficient condition for this is that (ii) is fulfilled [43, 44, 45] .
It is important to stress that of the two properties, (i) is definitely the most important, as it ensures that geometrically, the two spaces are properly matched. Instead, note that even if (ii) is not fulfilled, this will not jeopardize the stability of the learned policy. After transportation, data can be fit with a stable dynamical system, no matter the choice of $\phi$. Actually, the fact that we do not need to compute $J_{\phi}^{-1}$ is a key benefit of our method. As such, we will impose property (i) as part of the learning process, while property (ii) will only be checked experimentally a posteriori.
III-D A convenient choice for the transportation map
Taking inspiration from the definition of registration maps in the point cloud registration literature [42], we reformulate $\bm{\phi}$ as follows,
$\bm{\phi}(\bm{x}):=\bm{\gamma}(\bm{x})+\bm{\psi}(\bm{\gamma}(\bm{x})).$ (2)
Here, $\bm{\gamma}$ is an affine transformation and $\bm{\psi}$ a residual nonlinear transformation. This way, we ensure that the nonlinear transformation $\bm{\psi}$ only addresses the residual mismatch error remaining after the affine transformation $\bm{\gamma}$. The affine transformation will, therefore, act globally, and the nonlinearities will only locally add extra deformations to fine-tune the matching process. Fig. 2 shows an example of such a policy transformation in action.
Before diving into the proposed two-step process for learning $\bm{\phi}$ and $\bm{\psi}$, such that they fulfill property (i), we delve into some basic analytical considerations on conditions under which property (ii) is verified.
Proposition: If $\bm{\phi}$ fulfills property (i), then a necessary condition for it to also fulfill property (ii) is $\det(J_{\phi}(\bf{\mathcal{S}}))>0$ or $\det(J_{\phi}(\bf{\mathcal{S}}))<0$, where the inequality is to be intended element-wise. The condition becomes also sufficient if we ask it to be true across all possible choices of $\bf{\mathcal{S}}$.
Proof.
We wish to prove that $\det(J_{\phi})\neq 0$. The sufficient part is trivially implied by the hypothesis, as by taking all possible choices of $\bf{\mathcal{S}}$, we are sampling all the support of $J_{\phi}$.
For the necessary part, let us assume that the hypothesis is violated, i.e. there exist two elements $s_{i},s_{j}\in\bf{\mathcal{S}}$ such that $\det(J_{\phi}(s_{i}))<0$ and $\det(J_{\phi}(s_{j}))>0$. We define a generic continuous path $p:[0,1]\rightarrow\mathbb{R}^{3}$ such that $p(0)=s_{i}$ and $p(1)=s_{j}$. The function
$\det\circ J_{\phi}\circ p:[0,1]\rightarrow\mathbb{R}$ (3)
is continuous as it is a composition of continuous functions. Indeed, $\bm{\phi}$ is of class $C^{1}$ by hypothesis - i.e., its Jacobian is a continuous function. The determinant of a matrix is also a continuous function, and $p$ is a continuous by construction.
Thus, (3) is a continuous function starting negative in 0 and ending positive in 111. It follows that it must necessarily be zero at some point in $\bar{s}\in(0,1)$, thus confirming that violated hypothesis is necessary and concluding the proof.
III-E Learning Transportation Maps: $\phi$